Week of 160523

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes




  • local: Maria Dimou (SCOD), Maarten Litmaath (ALICE), Belinda Chan (CERN Storage), Julia Andreeva (WLCG Ops), Iain Steers (T0), Jose Andres Cordero (CERN DB), Marian Babik (networks), Marc Slater (LHCb)
  • remote: Francesco Noferini (CNAF), Dmytro Karpenko (NDGF), John Kelly (RAL), Alexander Verkooijen (NL_T1), Vincenzo Spinoso (EGI), Christoph Wissing (CMS), Rolf Rumler (IN2P3), Dario Barberis (ATLAS), Pepe Flix (PIC), Victor Zhiltsov (JINR), Dimitri (KIT), Kyle Gross (OSG).

Experiments round table:

  • ATLAS reports (raw view) -
    • Activities:
      • MC15c, derivation production, and Heavy Ion reprocessing ongoing - using 230k cores stably
      • ADC WM/DM review last week
    • Problems:
      • No particular problem this week apart from usual low-level data corruption and transfer troubles.

  • CMS reports (raw view) -
    • MC DigiReco for ICHEP still in full swing
    • Only minor operational issues, like this recent one

  • ALICE -
    • SAM: the critical profile for A/R computations now includes selected MonALISA metrics
      • see http://wlcg-sam-alice.cern.ch
      • in particular the SE services are now taken into account
      • the profile will be used as of May 1st (sic),
        but this first month we can be flexible with corrections
    • CERN: CREAM job submission incident on Sunday (GGUS:121699)
      • None of the CEs worked
      • Most had their sandbox partitions full
      • The main usage was due to ALICE jobs, apologies!
      • An ad-hoc cleanup was done
      • New jobs will by default not make use of output sandboxes
        • only needed for debugging
      • Job submissions resumed shortly before midnight

Maria asked whether this announcement of the new profile for a date that already passed is ok. Maarten replied that this was already announced at the ALICE workshop in Bergen in April. Marian verified that the support line remains the same: Nagios issues may be handled by the SAM team as usual, while tickets about MonALISA metrics will be forwarded to ALICE.

  • LHCb reports (raw view) -
    • Activity
      • Staging and Incremental stripping ~100% complete
    • Site Issues
      • T0: NTR
      • T1: NTR
Sites / Services round table:

  • ASGC: not connected
  • BNL: By Eric Lancon in email: NTR
  • CNAF:During last week, we completed the installation of additional 500 TB in the CMS storage which is now close to the pledged 2016.
  • FNAL: not connected
  • GridPP: not connected
  • IN2P3: NTR
  • KISTI: not connected
  • KIT: NTR
  • NDGF: https://its.cern.ch/jira/browse/ADCSUPPORT-4549 (but it's being looked at).
  • NL-T1: NTR
  • NRC-KI: not connected
  • OSG: NTR
  • PIC: The SIR on T10KD issues is ready here.
  • RAL:
    • There were problems with CMS storage access as in mentioned in GGUS :121698 - one disk server broken and all disk servers running out of file handles. We had believed file handle problem fixed on Friday but deployment system was not configured to restart services. Finally fixed this morning.
    • We are now deploying a new tranche of worker nodes, expect to have 3000 more job slots at RAL soon.
    • There was an issue with the RAL CVMFS stratum-1. Catalin was asked to provide an explanation and I have attached his email.
To give a summary to the WLCG ops meeting today about the last week problem on the RAL Stratum-1:

Last week between Monday (~late afternoon) and Wednesday (~noon) there were problems on the Stratum-1@RAL. Both at replication (cern.ch repositories) and client access levels because of some routing issues.

Explanation: There were some disk space issues on the production Stratum-1 HA cluster, and a standby stratum-1 (which took over for that period of time) didn't have the correct routing in place. Disk space issue resolved and the production HA stratum-1 servers were back online last Wednesday afternoon.

Please ask if more details needed.

  • TRIUMF: Nothing to report (will not join due to public holiday)

  • CERN computing services:
    • VOMS issues; Old node due to be retired accidentally re-entered the lb alias with a bad certificate from Tuesday (16th) afternoon until Thursday (18th) afternoon.
      • Unclear why the node had a certificate for the other voms service. This will be looked into by the Service Manager.
    • CREAM Sandbox Weekend issues; Series of ALICE jobs managed to fill the output sandboxes on the CREAM CEs, was manually cleaned up.
    • On-going CREAM Scalability/Stability issues: Ramp up in job submissions today (23rd) has resulted in sporadic but fairly common time-outs across the CREAM cluster.
  • CERN storage services: Nothing to report.
  • CERN databases: Nothing to report.
  • Network:
    • GGUS:119820 ASGC connectivity - tests performed didn't confirm Cisco N7K to be the only issue
    • GGUS:121687 RAL consistent loss - waiting for an upgrade of the router at RAL
  • GGUS:
  • MW Officer:
  • Security:


Topic attachments
I Attachment History Action Size Date Who Comment
Unknown file formatpptx GGUS-for-MB-May-16.pptx r1 manage 2817.0 K 2016-05-23 - 10:55 MariaDimou GGUS Slide for the 24 May MB
Edit | Attach | Watch | Print version | History: r18 < r17 < r16 < r15 < r14 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r18 - 2016-05-23 - MaartenLitmaath
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback